# Hints of Entanglement Suppression in Hyperon-Nucleon Scattering

### v1

@@ I Introduction
Understanding the dynamics of baryons is essential for obtaining a comprehensive picture of strong interactions, with broad implications in particle, nuclear, and astrophysics. While nucleon-nucleon interactions have been measured and constrained very precisely based on more than 4300 data points Stoks et al. (1993); Wiringa et al. (1995), interactions involving baryons carrying strangeness quantum number – the hyperon – are much less understood, due to experimental difficulties associated with the short lifetime of hyperon and the challenges of preparing a stable hyperon beam. Very little experimental progress was made on $YN$ scattering since the 1970s, until recently when the E40 collaboration at J-PARC and the CLAS collaboration at CEBAF announced new measurements Miwa et al. (2021, 2022); Nanamura et al. (2022); Rowley et al. (2021). Furthermore, there are also relevant measurements on $YN$ interactions at the Large Hadron Collider from the ALICE collaboration recently Acharya et al. (2020, 2022).
@@ II Kinematics
In the experimental setup, a beam of hyperon particles hit a stationary target, the nucleon. We label the incident particle mass as $m_{1}$ and the stationary particle mass $m_{2}$. Masses of outgoing particles are represented by $m_{1}^{\prime}$ and $m_{2}^{\prime}$. Experimental observables are usually measured as a function of the kinematic energy $T_{lab}$ and the magnitude of the 3-momentum ${P}_{lab}$ of the incident particle in the laboratory (Lab) frame. We will focus on the kinematic regime below the pion production threshold,
$N_{1}+N_{2}\to N_{1}^{\prime}+N_{2}^{\prime}+\pi\ ,$ (1)
so that the interaction is described by the pionless effective field theory Bedaque and van Kolck (1998); Kaplan et al. (1998a, b); van Kolck (1999). It is the easiest to calculate the pion threshold in the centre-of-mass (CM) frame which in terms of $p_{CM}$, the magnitude of the 3-momentum of the incident particle, is
@@ IV Global Fits
[…]
@@
[…]
There are two features of $\chi$EFT which are distinctly different from the global fits utilizing the potential models such as NSC97 and ESC16. First, $\chi$EFT is an expansion of the potential in small momenta and pion masses, augmented with an appropriate power counting rule. As such there is an inherent cutoff of $\chi$EFT, which is usually taken to be $\Lambda\sim 500$ MeV. This implies that, as the $p_{CM}$ gets close to $\Lambda$, higher order effects become important and $\chi$EFT starts breaking down. Second, $\chi$EFT employs the effective range expansion (ERE) Bethe (1949):
$p_{CM}\,\cot\delta=-\frac{1}{a}+\frac{1}{2}\Lambda^{2}\sum_{n=0}^{\infty}r_{n}\left(\frac{p_{CM}^{2}}{\Lambda^{2}}\right)^{n}.$ (11)
[…]
@@ V Results
Given that, in the case of $\Sigma^{+}p$, different global fits give rise to inconclusive outcome, we further propose “quantum observables” which could not only distinguish the varying global fits but also help determine the density matrix of the outgoing states Liu and Low . The observables are based on the formalism introduced in Ref. Hoshizaki (1969), which relates the density matrix of outgoing states to their polarizations. (For quantum observables in top quark decays, see Refs. Afik and de Nova (2021); Aad et al. (2023).) In Fig. 4 we plot the predicted polarizations, as a function of $p_{CM}$, of the recoiling hyperon (recoil polarization) and the recoiling proton (target polarization) in $\Sigma^{+}p$ scattering from NSC97 and ESC16 fits, assuming an unpolarized proton target and a 25% polarized incoming hyperon beam. We see that, by measuring the recoil and target polarizations, it is possible to distinguish between the two global fits. At J-PARC the $\Sigma^{+}$ particles come from the process $\pi^{+}p\to\Sigma^{+}K^{+}$ and is partially polarized in the order of 25%, depending on the incoming momenta Miwa and Nanamura . A more detailed study of such a scenario will be presented elsewhere Liu and Low .
@@ VI Conclusion
Hyperon-nucleon interactions are important for resolving the “hyperon puzzle,” which pertains to the formation of neutron stars heavier than two solar-masses. Inspired by recent experimental efforts in direct measurements of $YN$ scattering, as well as by theoretical advances in understanding nuclear dynamics from the perspective of quantum entanglement, we studied in this work the information-theoretic properties of two-body $YN$ scatterings, focusing on the question of whether the spin entanglement is suppressed during the scattering process. Using globals fits of scattering data, we find hints of entanglement suppression in the majority of $YN$ scattering channels. In the case of $\Sigma^{+}p$ scattering, conflicting global fits lead to inconclusive outcomes.

### v2

@@ I Introduction
Understanding the dynamics of baryons is essential for obtaining a comprehensive picture of strong interactions, with broad implications in particle, nuclear, and astrophysics. While nucleon-nucleon interactions have been measured and constrained very precisely based on more than 6700 data points Stoks et al. (1993); Wiringa et al. (1995); Gross and Stadler (2008); Machleidt (2001); Pérez et al. (2013), interactions involving baryons carrying strangeness quantum number – the hyperon – are much less understood, due to experimental difficulties associated with the short lifetime of hyperon and the challenges of preparing a stable hyperon beam. Very little experimental progress was made on $YN$ scattering since the 1970s, until recently when the E40 collaboration at J-PARC, the CLAS collaboration at CEBAF and the BESIII collaboration at BEPCII announced new measurements Miwa et al. (2021, 2022); Nanamura et al. (2022); Rowley et al. (2021); Ablikim et al. (2010, 2023). Furthermore, there are also relevant measurements on $YN$ interactions at the Large Hadron Collider from the ALICE collaboration recently Acharya et al. (2020, 2022).
@@ II Kinematics
In the experimental setup, a beam of hyperon particles hit a stationary target, the nucleon. We label the incident particle mass as $m_{1}$ and the stationary particle mass $m_{2}$. Masses of outgoing particles are represented by $m_{1}^{\prime}$ and $m_{2}^{\prime}$. Experimental observables are usually measured as a function of the kinematic energy $T_{lab}$ and the magnitude of the 3-momentum ${P}_{lab}$ of the incident particle in the laboratory (Lab) frame. We center on 2-to-2 scattering and therefore focus on the kinematic regime below the pion production threshold,
$N_{1}+N_{2}\to N_{1}^{\prime}+N_{2}^{\prime}+\pi\ .$ (1)
It is the easiest to calculate the pion threshold in the centre-of-mass (CM) frame which in terms of $p_{CM}$, the magnitude of the 3-momentum of the incident particle, is
@@ IV Global Fits
Figure 1: Comparisons of phase shifts in the $\Sigma^{+}p$ channel from potential models and N2LO $\chi$EFT ($\Lambda=$500 MeV).
@@
The second type of fits makes use of $\chi$EFT, which is pioneered by Weinberg Weinberg (1990, 1991) and has been very successful in describing $NN$ interactions. Applications of $\chi$EFT to $YN$ interactions have progressed steadily over the years Polinder et al. (2006); Haidenbauer et al. (2013, 2020) and the state-of-the-art calculation now stands at next-to-next-to-leading order (N2LO) Haidenbauer et al. (2023), which we use.
There is one feature of $\chi$EFT which is distinctly different from the global fits utilizing the potential models such as NSC97 and ESC16. $\chi$EFT is an expansion of the potential in small momenta and pion masses, augmented with an appropriate power counting rule. As such there is an inherent cutoff of $\chi$EFT, which is usually taken to be $\Lambda\sim 500$ MeV. This implies that, as the $p_{CM}$ gets close to $\Lambda$, higher order effects become important and $\chi$EFT starts breaking down. In Fig. 1 we show a comparison of the phase shifts in the $\Sigma^{+}p$ channel from the potential models and from $\chi$EFT. We see that phase shifts from ESC16 and N2LO are consistent up to $p_{CM}\sim 250$ MeV. Moreover, the ${}^{3}S_{1}$ phase shifts in ESC16 and N2LO $\chi$EFT have a negative sign (repulsive force) while in NSC97 it has a positive sign (attractive force). In the isospin-related channel of $\Sigma^{-}n$, lattice simulations seem to also indicated a negative ${}^{3}S_{1}$ phase shift Beane et al. (2012).
@@ V Results
Given that, in the case of $\Sigma^{+}p$, different global fits give rise to inconclusive outcome, we further propose “quantum observables” which could not only distinguish the varying global fits but also help determine the density matrix of the outgoing states Liu and Low . The observables are based on the formalism introduced in Ref. Hoshizaki (1969), which relates the density matrix of outgoing states to their polarizations. (For quantum observables in top quark decays, see Refs. Afik and de Nova (2021); Aad et al. (2023).) In Fig. 4 we plot the predicted polarizations, as a function of $p_{CM}$, of the recoiling hyperon (recoil polarization) and the recoiling proton (target polarization) in $\Sigma^{+}p$ scattering from NSC97 and ESC16 fits, assuming an unpolarized proton target and a 25% polarized incoming hyperon beam. We include phase shifts up to d wave as well as the ${}^{3}S_{1}-^{3}D_{1}$ mixing. We see that, by measuring the recoil and target polarizations, it is possible to distinguish between the two global fits. At J-PARC the $\Sigma^{+}$ particles come from the process $\pi^{+}p\to\Sigma^{+}K^{+}$ and is partially polarized in the order of 25%, depending on the incoming momenta Miwa and Nanamura . A more detailed study of such a scenario will be presented elsewhere Liu and Low .
@@ VI Conclusion
Hyperon-nucleon interactions are important for resolving the “hyperon puzzle,” which pertains to the formation of neutron stars heavier than two solar-masses. Inspired by recent experimental efforts in direct measurements of $YN$ scattering, as well as by theoretical advances in understanding nuclear dynamics from the perspective of quantum entanglement, we studied in this work the information-theoretic properties of two-body $YN$ scatterings, focusing on the question of whether the spin entanglement is suppressed during the scattering process. Using globals fits of scattering data, we find hints of entanglement suppression in the majority of $YN$ scattering channels. In the case of $\Sigma^{+}p$ scattering, only the more dated NSC97 fit showed entanglement suppression, while the more recent fits do not demonstrate entanglement suppression in the channel.